### 09 Feb 2024 v1

# Value function interference and greedy action selection in value-based multi-objective reinforcement learning

Abstract
Multi-objective reinforcement learning (MORL) algorithms extend conventional reinforcement learning (RL) to the more general case of problems with multiple, conflicting objectives, represented by vector-valued rewards. Widely-used scalar RL methods such as Q-learning can be modified to handle multiple objectives by (1) learning vector-valued value functions, and (2) performing action selection using a scalarisation or ordering operator which reflects the user’s utility with respect to the different objectives. However, as we demonstrate here, if the user’s utility function maps widely varying vector-values to similar levels of utility, this can lead to interference in the value-function learned by the agent, leading to convergence to sub-optimal policies. This will be most prevalent in stochastic environments when optimising for the Expected Scalarised Return criterion, but we present a simple example showing that interference can also arise in deterministic environments. We demonstrate empirically that avoiding the use of random tie-breaking when identifying greedy actions can ameliorate, but not fully overcome, the problems caused by value function interference.
Keywords multi-objective reinforcement learning  $\cdot$ Q-learning  $\cdot$ value-based reinforcement learning  $\cdot$ action selection
1 Introduction and Background
Multi-objective reinforcement learning (MORL) is a form of reinforcement learning (RL) designed for problems with multiple, conflicting objectives (Roijers et al., 2013). Whereas conventional, single-objective RL algorithms assume that the agent is provided with a scalar reward after each action, in MORL these rewards are vectors with a separate component for each objective. Prior research has shown that applying single-objective methods to multiobjective tasks may result in unsuitable, inferior or inflexible solutions (Roijers et al., 2013, 2015), and so in recent years there has been considerable interest in developing MORL methods. In many cases these MORL algorithms are extensions of existing RL methods, including value-based approaches such as Q-learning (Gábor et al., 1998; Natarajan and Tadepalli, 2005; Van Moffaert et al., 2013; Van Moffaert and Nowé, 2014; Ruiz-Montiel et al., 2017; Vamplew et al., 2017a).
Adapting value-based reinforcement learning methods to work with multiple objectives requires two key changes to be made:
  1. 1.
    As the rewards received by the agent are vectors, the Q-values stored by the agent must also be vectors.
  2. 2.
    When action selection is performed, a greedy action must be identified in a way which is consistent with the user’s overall utility Roijers et al. (2013); Zintgraf et al. (2015).
The action selection can in most cases be implemented via applying a scalarisation operator which maps the vector-values to scalars, and then regarding the action with the highest scalarised value as being the greedy selection. Depending on the user’s utility preferences this might be a simple linear weighting or a non-linear operator, such as the Chebyshev distance [^1] We note that when a non-linear operator is used, then both the action selection and the Q-values learned by the agent must be conditioned not just on the current state of the environment, but also on some measure of the rewards received so far (Vamplew et al., 2021). An exception is in the case where rewards are zero except at terminal states (Issabekov and Vamplew, 2012). As this issue is orthogonal to those considered in this paper, for clarity we will only consider this latter form of environment within this paper - however the observations made here are equally applicable to the more general case.. For some forms of utility (e.g. those based on lexicographic ordering, which often arise in the context of meeting constraints (Gábor et al., 1998) or in the problems where fairness is a consideration (Siddique et al., 2020)), it may not be possible to represent the utility directly in terms of scalarisation (Debreu, 1997). In such cases, the Q-learning agent can instead use an ordering operator to identify the greedy policy.
Algorithm 1 provides a generic form of multi-objective Q-learning. The action-selection function $U$ may be based either on explicit scalarisation of value-vectors or on an ordering operator, depending on the nature of the user’s utility function.
Algorithm 1 A general algorithm for multi-objective Q($\lambda$).
1:input: learning rate $\alpha$, discounting term $\gamma$, eligibility trace decay term $\lambda$, number of objectives $n$, action-selection function $U$ and any associated parameters
2:for all states $s$, actions $a$ and objectives $o$ do
3:     initialise $Q_{o}(s,a)$
4:end for
5:for each episode do
6:     for all states $s$ and actions $a$ do
7:         $e(s,a)$=0
8:     end for
9:     sums of prior rewards $P_{o}$ = 0, for all $o$ in 1..$n$
10:     observe initial state $s_{t}$
11:     $s_{t}$ = $(s_{t},P)$triangleright create augmented state
12:      select $a_{t}$ from an exploratory policy derived using $f(Q(s))$
13:     for each step of the episode do
14:         execute $a_{t}$, observe $s_{t+1}$ and reward $R_{t}$
15:         $P=P+R_{t}$
16:         $s_{t+1}$ = $(s_{t+1},P)$triangleright create augmented state
17:         $V(s_{t+1})=Q(s_{t+1})+P$triangleright create value vector
18:         select $a^{*}$ from a greedy policy derived using $U(V(s_{t+1}))$
19:         select $a^{\prime}$ from an exploratory policy derived using $U(V(s_{t+1}))$
20:         $\delta=R_{t}+\gamma Q(s_{t+1},a^{*})-Q(s_{t},a_{t})$
21:         $e(s_{t},a_{t})$ = 1
22:         for each state $s$ and action $a$ do
23:              $Q(s,a)=Q(s,a)+\alpha\delta e(s,a)$
24:              if $a^{\prime}=a^{*}$ then
25:                  $e(s,a)=\gamma\lambda e(s,a)$
26:              end if
27:         end for
28:         $s_{t}=s_{t+1},a_{t}=a^{\prime}$
29:     end for
30:end for
One further consideration in MORL is the choice of optimisation criteria, as there are two distinct options compared with just a single possible criteria in conventional RL (Roijers et al., 2013). The first one is to maximise the Expected Scalarised Return (ESR). In this approach, the agent aims to maximise the expected value which is first scalarised by the utility function, as shown below (Eq 1) where $w$ is the parameter vector for utility function $f$, $r_{k}$ is the vector reward on time-step $k$, and $\gamma$ is the discounting factor
$V^{\pi}_{\bf w}(s)=E[f(\sum^{\infty}_{k=0}\gamma^{k}\mathbf{r}_{k},{\bf w})\ |\ \pi,s_{0}=s)$ (1)
ESR is the appropriate criteria for problems where we care about the outcome within each individual episode. For example, consider selecting a treatment plan for a patient, where there is a trade-off between the effectiveness of the treatment and negative side-effects. Each patient would care about their own individual outcome rather than the mean outcome over all patients.
The second criteria is the Scalarised Expected Return (SER) which estimates the expected vector return per episode and then maximises the scalarised expected return, as shown in (Eq 2)
$V^{\pi}_{\bf w}(s)=f({\bf V}^{\pi}(s),{\bf w})=f(E[\sum^{\infty}_{k=0}\gamma^{k}\mathbf{r}_{k}\ |\ \pi,s_{0}=s],{\bf w})$ (2)
SER is an appropriate criteria when we care about the optimal utility over multiple executions of a policy. For example a logistics company might seek to optimise both delivery time and fuel consumption, but as they make many deliveries they care about the average performance across all episodes rather than the outcome from each individual episode.
Multiobjective extensions of Q-learning have been amongst the most widely used forms of MORL reported so far in the literature (Gábor et al., 1998; Natarajan and Tadepalli, 2005; Van Moffaert et al., 2013; Van Moffaert and Nowé, 2014; Ruiz-Montiel et al., 2017; Vamplew et al., 2017a). However, as we will demonstrate via example and empirically in this paper, when used in combination with a non-linear utility function an issue which we will label value function interference can arise, which may slow learning or cause the agent to converge to a sub-optimal policy.
2 Interference in multi-objective value functions for deterministic environments
We start by considering the simplest case, which are environments where the choice of starting state, the state transitions and the rewards are fully deterministic. Assuming that the policy being executed is also deterministic (as would usually be the case for methods based on extensions of Q-learning), then in this context the SER and ESR optimisation criteria are equivalent (every execution of a policy results in the same return, which of course is also the same as the mean return considered by SER).
2.1 An example of value function interference in a stochastic MOMDP
As an illustrative example, consider the simple deterministic multi-objective Markov Decision Process (MOMDP) shown in Figure 1. This MOMDP consists of three states ($A$, $B$, and $C$). In each state two actions are available ($a_{1}$ and $a_{2}$). The MOMDP has three objectives, so the reward vector has three elements. There are four possible policies, as summarised in Table 1.
Figure 1: An example multi-objective MDP. The unlabelled states are terminal states, and the vector reward for reaching these states is indicated to their right. All other state transitions receive zero rewards.
Policy label Action in State A Action in State B Action in State C Vector return Utility
0 $a_{1}$ $a_{1}$ (7, -1, -5) 9
1 $a_{1}$ $a_{2}$ (7, -5, -1) 9
2 $a_{2}$ $a_{1}$ (8, -3, -3) 7
3 $a_{2}$ $a_{2}$ (0, -5, -5) -25
Table 1: The four policies for the MOMDP from Figure 1. The scalar utility in the final column assumes that the utility equals twice the first objective minus the product of the second and third objectives.
There are many ways in which a utility function could be derived from the rewards [^2] The actual specification of rewards and utilities is outside the scope of this project; we focus here simply on examining whether MORL methods will optimise for the utilities and rewards as defined by the user, irrespective of the structure of those definitions. The issues identified here could arise for any non-linear definition of utility., but let us assume that $U$ equals twice the value of the first objective minus the product of the second and third objectives (which are always negative). Under these settings $U(B,a_{1})=U(B,a_{2})=9,U(C,a_{1})=7$, and $U(C,a_{2})=-25$. Therefore clearly the optimal policy should be to select action $a_{1}$ in state $A$, followed by either action in state $B$ i.e. either policy 0 or policy 1 (this is true for both the SER and ESR criterion).
However the existence of two equally optimal actions in state $B$ creates an issue in the context of multi-objective rewards. The usual approach taken in Q-learning where more than one optimal action exists is to simply randomly select between them. However this random selection introduces an element of stochasticity into the vector rewards which the agent receives, which in turn leads to interference in the values which the agent learns. This interference problem does not arise in single-objective RL, because two actions which are equally optimal must, by definition, share the same mean scalar reward, whereas in the multi-objective case two actions may share the same utility while having differing mean vector rewards.
Consider the values learned for $Q(A,a_{1})$. This action always leads to state $B$, but at that stage under the optimal policy either action is equally likely to be selected, and so $Q(A,a_{1})$ will converge to the mean of $Q(B,a_{1})$ and $Q(B,a_{2})=(7,-3,-3)$, which has an estimated utility of just 5. Meanwhile because action $a_{2}$ in state $A$ will always be followed by action $a_{1}$ in state $C$, $Q(A,a_{2})$ will converge to the same value as $Q(C,a_{1})$, and so will have utility of 7. Therefore ultimately the agent will converge incorrectly to the sub-optimal policy 2 (selecting action $a_{2}$ in state $A$ and $a_{1}$ in state $C$), with a considerable consequent loss of utility.
This example illustrates the core concept of value function interference in value-based multi-objective RL. If the utility function allows actions with widely differing vector outcomes to map to the same utility value, then the vector value function learned for earlier states may be inconsistent with the actual optimal policy. We note that if the utility function is linear then value interference does not impact on selecting the optimal action. The value learned for $Q(A,a_{1})$ will lie (approximately) on the line segment joining between $Q(B,a_{1})$ and $Q(B,a_{2})$. If $Q(B,a_{1})$ and $Q(B,a_{2})$ share the same utility value for a particular linear weighting of objectives, then all points on that line also have that same utility.
2.2 Empirical evaluation of a potential solution
The problem of value function interference in a stochastic MOMDP can potentially be addressed by modifying the selection process so that ties between multiple greedy actions are broken in a deterministic fashion. For example, we might always select the greedy action with the lowest sub-script - this introduces some degree of bias in the action selection, but presumably the user is indifferent to this bias (otherwise it would be represented in some way within the utility function).
To demonstrate the impact of value function interference, and the ability of non-random tie-breaking to mitigate its effects, empirical experiments were run on the MDP from Figure 1. These experiments used three variations of a tabular multi-objective Q-learning agent. The first variant is a baseline using the standard random approach to tie-breaking in the case of actions with equal scalarised values. The second variant deterministically breaks ties by favouring the action with the lower index (i.e. $a_{1}$ is preferred to $a_{2}$). As this could produce a bias towards the desired policies (both of which select $a_{1}$ in the first state), we also included a third variant which favours the action with the higher index.
Each agent’s performance was evaluated over a hyperparameter sweep across different values for the learning rate ($\alpha$) and starting exploration parameter (using e-greedy exploration with $\epsilon$ linearly decayed to zero over the trial). The $\lambda$ and $\gamma$ hyperparameters were fixed at 0.95 and 1 respectively, and Q-values were optimistically initialised to (12,0,0). For each set of hyperparameters 100 trials of each agent were executed. Each trial lasted for 500 episodes, and at the end of the trial the final greedy policy of the agent was identified.
Random tie-breaks
Low-index tie-breaks
High-index tie-breaks
Figure 2: Heatmaps showing the frequency with which each policy was selected as the final greedy policy across 100 trials using different hyperparameter settings for each of the three different tie-breaking approaches: random (top row), lower-indexed action (middle row), and higher-indexed action (bottom row)
Figure 2 is a heatmap showing the frequency of occurence of each of the possible policies for each agent/hyperparameter combination. Policy 3 has been omitted as it was never selected as the final policy in any trial.
2.2.1 Random tie-breaking
The results from the random tie-breaking agent (Figure 2, top row) clearly illustrate the problems caused by value function interference, as for large areas of the hyperparameter sweep the agents converged to the sub-optimal Policy 2 in the overwhelming majority of trials. Notably this was most likely to occur for trials where the learning rate $\alpha$ was in the middle of its range $0..1$.
This can be explained by the nature of the utility function. The value learned for $Q(A,a_{1})$ will over time converge to lie on the line segment between (7,-1,-5) and (7, -5, -1) (i.e. between the rewards for $B,a_{1}$ and $B,a_{2}$. The location along this line will depend on the frequency with which each action has been performed in state $B$ recently, as well as the magnitude of $\alpha$. Points on this line-segment will have the form $p=(7,-5+4x,-1-4x)$ where $0\leq x\leq 1$. Action $a_{1}$ will be preferred to $a_{2}$ in state $A$ when $U(Q(A,a_{1}))-U(Q(A,a_{2}))>0$. Substituting the definition of $p$ and the value of $Q(A,a_{2})$ yields that this will occur when $016x^{2}-16x+2>$ 0, which occurs when $x<(2-\sqrt{(}2))/4$ and when $x>(2+\sqrt{(}2))/4$ (i.e. when $Q(A,a_{1})$ is close to either $Q(B,a_{1})$ or $Q(B,a_{2})$.
Consider the situation when $Q(A,a_{1})$ is a good approximation to $Q(B,a_{1})$, and now action $B,a_{2}$ is selected as a greedy action and executed. $Q(A,a_{1})$ will be updated towards (7, -5, -1). A small value of $\alpha$ will produce a small change in $Q(A,a_{1})$ meaning it will likely still be in the range where it is considered preferable to $Q(A,a_{2})$. A large value of $\alpha$ will produce a large change in $Q(A,a_{1})$, jumping over the portion of the line-segment where $A,a_{2}$ is preferred. Meanwhile middling values of $\alpha$ will result in $Q(A,a_{1})$ falling roughly halfway between (7,-1,-5) and (7, -5, -1) (i.e. $x\approx 0.5$), which is the region where $A,a_{2}$ will be preferred. Hence it is this central range of $\alpha$ values where convergence to the incorrect policy is most likely to occur, which is evident in the experimental results.
There is also a tendency for sub-optimal convergence to become more frequent across a wider range of $\alpha$ values as the degree of exploration is increased. We note that the exploration values used here may seem high for such a simple MOMDP, but past research has shown that exploration rates may need to be higher for MORL than for single-objective RL (Vamplew et al., 2017b). Therefore this link between value function interference and high rates of exploration may prove to be problematic for more complex MOMDPs.
2.2.2 Deterministic tie-breaking
The results from the two deterministic tie-breaking agents (Figure 2, middle and bottom rows) demonstrate substantial improvements over the random tie-breaking agent (most notably for the case where $\alpha$=0.7, but more generally across the central portions of the hyperparameter search). Clearly the avoidance of random tie-breaking helps to mitigate, but not entirely avoid, the problems caused by value function interference – there are still combinations of hyperparameters where most/all trials converge to the incorrect final policy.
Comparing the results of the two deterministic tie-breaking agents, we see a clear effect of the process used to break ties – the first agent favours actions with a lower index, and so when it converges to one of the optimal policies, it has a clear preference for Policy 0 over Policy 1, whereas the second agent favours actions with a higher index and exhibits the opposite preference over these policies.
Given this capacity for the deterministic tie-breaking to induce different preferences over policies, we need to consider the possibility that the reduction in the frequency of convergence to sub-optimal policies is not due directly to the elimination of random tie-breaking, but instead due to an explicit bias away from Policy 2. Figure 3 assists in differentiating between these causes by performing a pairwise comparison of the agents based on how frequently they incorrectly converge to Policy 2. It can be seen from Figure 3(left) that the low-index deterministic tie-breaking agent converges to this policy with lower or comparable frequency to the random tie-breaking agent for every hyperparameter setting which was tested. Overall it selects this policy in 1493 fewer trials than the random tie-breaking agent (almost 30% of the total number of trials). The high-index deterministic tie-breaking agent also improves substantially over the random tie-breaking agent, selecting the sub-optimal policy in 1385 fewer cases, but this is 108 cases worse than the low-index agent (Figure 3 (middle and right)). So while the low-index agent’s bias away from action $a_{2}$ in state $A$ is partially responsible for the reduction in trials converging to Policy 2, this accounts for just 108/1493 cases – more than 90% of the improvement is due to avoiding the stochasticity of the random tie-breaker rather than the specific bias of the deterministic tie-breaker.
Random tie-breaks
Figure 3: Heatmaps showing the difference in the frequency with which the incorrect Policy 2 was selected as the final greedy policy across 100 trials using different hyperparameter settings for each combination of the three different tie-breaking approaches: Higher-index -random (top row); Lower-indexed - random (middle row); lower-index - higher-indexed (bottom row)
These results demonstrate that deterministic tie-breaking may be beneficial in the case of MOMDPs where some actions share identical utility. However more generally the actions need not map to exactly the same scalar utility value in order for interference to arise. In practice the Q-value estimates may be noisy due to either the use of function approximation and/or stochasticity in the environment. These errors in estimates may result in changes in the action which is selected as greedy in each state, leading to variations over time in the vector value updates for prior states, which could result in value function interference. This is particularly likely to occur for highly non-linear definitions of utility as small variations in the estimated vector values can lead to large variations in the estimated utility; for example, we have previously observed unstable and slow learning due to value function interference when using a thresholded lexicographic approach to utility on stochastic problems (Vamplew et al., 2020, 2021; Ding et al., 2024).
3 Interference in multi-objective value functions for stochastic environments
If we now consider environments with some element of stochasticity (in either the starting states, state transitions or rewards), then the SER and ESR criteria are no longer equivalent, and value interference issues may arise for ESR-optimal learning which do not occur for SER-optimal learning.
Figure 4: An example multi-objective MDP with stochastic rewards. The unlabelled states are terminal states, and the vector reward for reaching these states is indicated to their right.
Consider the very simple MOMDP  [^3] In fact, as this MOMDP has only a single non-terminal state, this example demonstrates that value function interference can arise even for ESR optimisation of multi-objective multi-armed bandits (MOMABs). shown in Figure 4. Action $a_{1}$ has a stochastic reward which varies between (7, -1, -5) and (7, -5, -1) with equal probability. The utility of this action differs depending on whether we are optimising with respect to ESR or SER – assuming the same definition of utility as in Section 2, $U_{SER}(a_{1})$=5 (the scalarised value of the expected vector reward) while $U_{ESR}(a_{1})$=9 (the expected scalarised value, which is 9 for both of the possible outcomes). Action $a_{2}$ deterministically returns a reward of (8, -3, -3) – as this is deterministic $U_{SER}(a_{2})=U_{ESR}(a_{2})$ = 7.
Therefore the SER-optimal policy in this case is to select $a_{2}$, while the ESR-optimal policy selects $a_{1}$. The SER-optimal policy can be determined on the basis of the vector Q-values learned by the agent, but this is not true for the ESR-optimal policy, as these values only represent the expected vector return which is insufficient for the agent to determine the per-episode expected scalarised return. The use of deterministic tie-breaking proposed as a mitigating solution earlier in Section 2 is not useful in this context, as the stochasticity in the returns arises from the environment rather than being induced by stochasticity in the policy due to tie-breaking.
For this scenario we suggest that the most promising course of action is to incorporate concepts from distributional RL into MORL, as pioneered by Hayes et al. (2021); Reymond et al. (2023); Röpke et al. (2023). If the agent learns a distribution over the vector returns for each state-action, then this can be combined with the utility function so as to estimate the ESR. For example if the agent has correctly learned that the distribution of future returns for action $a_{1}$ is (-7,-1, -5) with $p=0.5$ and (-7,-5, -1) with $p=0.5$. From this information the agent can correctly estimate the ESR utility of $a_{1}$ as the probabilistically-weighted sum of those return vectors.
$\begin{split}U_{ESR}(a_{1})=0.5U_{ESR}(7,-1,-5)+0.5U_{ESR}(7,-5,-1)\\ =0.5*9+0.5*9\\ =9\end{split}$ (3)
4 Conclusion
Value function interference is an issue affecting value-based approaches to multi-objective reinforcement learning, which has not previously been identified in the MORL literature. These methods usually learn Q-values which are estimates of the expected mean future vector return, and in some cases this is not sufficient information to establish the optimal policy. This is particularly evident when trying to learn ESR-optimal policies for stochastic environments, but this paper has also demonstrated that it is possible for value function interference to arise in the context of SER-optimal policies for deterministic environments. We have shown empirically that in the latter case, modifying the core Q-learning algorithm to use non-random tie-breaking when two actions have equivalent utility can ameliorate, but not totally eliminate, the impact of value function interference. More generally, we advocate the use of alternative approaches such as distributional forms of value-based MORL or policy-search MORL methods to eliminate or avoid this issue.